OpenStudy (anonymous):

Find the surface area of a sphere that has a volume of 288 pi cu. in. ? pi

OpenStudy (anonymous):

???

OpenStudy (radar):

Do you have the formula for the surface area of a sphere? I can provide that if you want?

OpenStudy (radar):

Sphere surface area, SA=4pir^2

OpenStudy (radar):

Sphere Volume, V=(4/3)pi r^3

OpenStudy (anonymous):

4/3*pi*r^3=288pi r=6 then surface=4*pi*r^2=4*pi*36 =144pi

OpenStudy (radar):

Solve for r. from given volume solve for SA, ur done.

OpenStudy (radar):

Hey rathon has done it for you. good luck.

OpenStudy (anonymous):

i am in china

OpenStudy (radar):

Hi from the Ozarks in Missouri , USA where it is below freezing and a dusting of snow.

OpenStudy (radar):

Raython, I like your user name, it has a nice "electronic" connection.

OpenStudy (anonymous):

amazing , i love computer ,i am trying study assemble language programming

OpenStudy (anonymous):

we are very far away from each other

OpenStudy (radar):

That is a good area of study, I am retired electonic person from FAA (Federal Aviation Administration) .I worked at long range en route radars here in USA. Good luck in assembly and machine language.

OpenStudy (radar):

Yes we are far apart. I have been closer having served in Japan and South Korea.

OpenStudy (anonymous):

i appriciate people like you , a fan .

OpenStudy (radar):

Nice meeting you. JohnyBoy do you have any questions?

OpenStudy (anonymous):

nice to meet you too ,